In a certain population, the sensitivity of a diagnostic test is 0.80, and the specificity is 0.90.

The prevalence of the disease in the population is 0.10. Use this information to answer Ques-

tions 11 to 13.

11. What is the probability of testing negative?

(a) 0.17

(b) 0.20

(c) 0.47

(d)

0.83

12. Among those who have the disease of interest, what is the probability that they test

negative?

(a) 0.10

(b)

0.20

(c) 0.47

(d) 0.83

13. Among those who test positive, what is the probability that they have the disease?

(a) 0.10

(b) 0.20

(c)

0.47

(d) 0.8
Answer #1

from above:

have not have total
positve 0.08 0.09 0.17
negative 0.02 0.81 0.83
total 0.1 0.9 1

11)  probability of testing negative =0.1*(1-0.8)+(1-0.1)*0.9 =0.83

12) P(negative|disease) =1-0.8 =0.2

13) P(have disease given tested positive) =0.1*0.8/(0.1*0.8+0.9*0.1)=0.08/0.17 =0.47